A square graphed on the coordinate plane has a diagonal with endpoints E(2, 3) and F(0, -3).
What are the coordinates of the endpoints of the other diagonal?


A. (4,−1) and (−3,−1)
B. (4,−1) and (−2,1)
C. (4,0) and (−2,−1)
D. (3,−1) and (−2,−1)

Found 2 solutions by mathsolverplus, AnlytcPhil:
Answer by mathsolverplus(88)

The midpoint of the given pairs of coordinates should be equal to the midpoint
of the answered pairs.
Midpoint of EF =
((2+0)/2, (3-3)/0) = (1, 0)
A: midpoint = ((4-3)/2, (-1-1)/2) = (0.5, -1)
B: midpoint = ((4-2)/2, (-1+1)/2) = (1, 0) <----Answer
C: midpoint = ((4-2)/2, (-1+0)/2) = (1, -0.5)
D: midpoint = ((3-2)/2, (-1-1)/2) = (0.5, -1)


If this answer has helped you and you need help on additional problems, please follow us on Instagram or message us via social media! We also help with SAT Math Questions!

***Instagram (@mathsolver.plus): https://www.instagram.com/mathsolver.plus/
Twitter (@MathSolverPlus): https://twitter.com/MathSolverPlus
Snapchat (@mathsolverplus)
Email: mathsolverplus@gmail.com

Answer by AnlytcPhil(1806)



Draw a vertical rectangle with those points at opposite corners.



Now draw horizontal rectangle just like it to make a big "plus" sign:



Draw in the square:



Now you can see that they are (4,-1) and (-2,1)

Edwin